# Surface bubble lifetime in the presence of a turbulent air flow, and the effect of surface layer renewal

### v2

@@ 4.3 Lifetime of solutions of pure surfactants
Figure 6: (a) Mean lifetime as a function of the inverse of the magnitude of the evaporative rate. (b) Dimensionless lifetime as a function of the time $t_{c}$ (defined in Eq. (10)) when evaporation becomes the main draining mechanism. Markers show relative humidity and colors show SDS concentration. (orange line) Numerical integration of Eq. (5), with $\alpha_{d}=0.5$ and $\alpha_{e}=12.5$. (black dashed line) Scaling law $t_{f}=0.4t_{c}$.
@@
In this section, we analyze the lifetime and drainage of surface bubbles in a solution only containing SDS (no salts). Plotting our data for several wind speeds, relative humidities and SDS concentrations (Fig. 6 (a)) reveals that most of the changes in mean lifetime are captured using a single parameter that contains the effects of wind and humidity on the evaporative rate: $\left(Sh\times(1-\mathcal{R}_{H})\right)^{-1}$. This result confirms that our modelling of the Sherwood number in the large range of conditions (from quiescent to fully turbulent) is correct. In addition, it shows that there is no visible effect of the wind on bubble lifetime beyond altering the evaporative rate. In this case with no salts in the solution, the effects of the surfactant concentration are very small and do not show a significant trend. Under the conditions explored in this study, the mean lifetime of bubbles in solutions containing only deionized water and surfactants is always larger than about 10 seconds (even though individual bubbles still sometimes burst earlier than 10s, see Fig 4 (b)).
We expect the lifetime of these bubbles to be of the same order of magnitude as the time $t_{c}$ defined in Eq. 10, when evaporation becomes the main draining mechanism. To test our scaling, we plot the same lifetime data as a function of $t_{c}$ (both times normalized by $\tau_{d}$) in Fig. 6 (b). This rescaling indeed captures well the evolution of the lifetime as $t_{f}/\tau_{d}$ grows linearly with $t_{c}/\tau_{d}$ with a slope of 0.4 (dashed line). Most of the spread in the data at high wind speed (lower left corner) that was present in Fig. 6 (a) is also captured by using $t_{c}/\tau_{d}$ showing that it was most likely due to small variations in surface tension or bubble size.
The same result can also be obtained numerically: we set a fixed thickness bursting criterion at a very small thickness ($h_{c}=10^{-3}h_{0}$), integrate Eq. (5) until the thickness reaches $h_{c}$ and record the associated time. This allows us to obtain the orange line, that also follows $t_{f}\sim t_{c}$ (shown analytically if $h_{c}=0$ in (Poulain and Bourouiba,, 2018)) and predicts the data very well if the prefactor $\alpha_{e}$ is slightly adjusted compared to its value obtained in the drainage experiments (here $\alpha_{d}=0.5$ and $\alpha_{e}=12.5$). This change is most likely due to small experimental differences between the drainage and lifetime setups: the tank in which the drainage experiment is performed is smaller such that the bubble always remains in view of the high speed camera. The evaporative prefactor $\alpha_{e}$ being larger in that case can therefore be interpreted as the influence of the proximity of the bath on the evaporative rate being less important.
@@
Poulain and Bourouiba, (2018) observed for bubbles in the large surfactant concentration limit, that the longest time a surface bubble can live is given by $t_{c}$, when evaporation becomes the dominant drainage mechanism. With our setup in which we took great care in removing impurities (see Sec 3), we find that the average lifetime is of order $t_{c}$ (instead of just the maximum). This result is valid for very large changes of SDS concentration, from values greater than the CMC down to just a few µmol/L (or $10^{-3}$ times the CMC). All of our lifetime data for SDS is well collapsed when plotted as a function of $t_{c}$. This result is particularly robust at high wind speeds (or low $t_{c}$), which is coherent with our previous observations that mixing the air around the bath and the bubble improves the repeatability of lifetime experiment by preventing local fluctuations of the humidity field. At high $t_{c}$, we observe lifetimes that can be larger than the trend $t_{f}/\tau_{d}=0.4\>t_{c}/\tau_{d}$, and in general more important fluctuations. These may be due to the difficulty in modeling the Sherwood number associated with natural convection.
@@ 4.4 Lifetime of solutions containing salts
The lifetime of bubbles in solutions of salt and surfactants is drastically different from that of bubbles in pure water and surfactants. The lifetime data is summed up in Fig. 7 for sea salt (a mixture of salts and inorganic compounds in their oceanic proportions) and NaCl only (35 g/L). The top row shows the dimensional lifetime as a function of the inverse of the magnitude of the evaporative rate $\left(Sh\times(1-\mathcal{R}_{H})\right)^{-1}$, illustrating that in this case too, this parameter is adequate to combine the effects of changing wind speeds and relative humidities. The bottom row shows the same data, made dimensionless in the same way as in Fig. 6, using the drainage timescale $\tau_{d}$. In contrast to the previous case where concentration did not have a large effect and all bubble burst at a time given by $t_{c}$, here we see in both cases a transition between a regime where bubble lifetime is much shorter than $t_{c}$ and independent of relative humidity or wind speed, and a regime again dominated by evaporation, with the lifetime given by Eq. (10), with the same prefactor as in Fig. 6 for pure water and SDS.
@@
At low concentrations ($c_{\textsc{SDS}}\lesssim 50$ µmol/L), the lifetime is essentially independent of wind speed or relative humidity. In this regime, the lifetime of solutions with sea salt is about 5 s and 10 s with NaCl. It is interesting to note that in both cases in this limit the lifetime is again independent on the surfactant concentration within the range tested. Even though the concentrations in SDS are fairly small (less than 1% of the CMC), the combined presence of salts and surfactants results in a large variation of the surface tension (see Appendix A).
@@
Figure 7: (a) and (b) Lifetime as function of inverse of the magnitude of the evaporative rate for the sea salt and NaCl cases, respectively. (c) and (d) Dimensionless lifetime as a function of the dimensionless time when evaporation becomes dominant $t_{c}/\tau_{d}$. The left column shows the dataset with sea salt and the right one with NaCl. The dashed lines are identical to Fig. 6 (c): linear trends with a slope of 0.4. The orange lines are obtained by integrating of Eq. (5) until the thickness reaches $h_{c}$ (with the same prefactors as in Fig. 6). Markers show relative humidity and color shows SDS concentration.
@@
Increasing the SDS concentration from this limit, we see the lifetime in both cases start to increase and to show a positive trend with the timescale $t_{c}$. In intermediate cases (100 and 200 µmol/L) the increase of lifetime with $t_{c}$ is slower than predicted by Eq. (10) and the lifetime always remains inferior to the pure water and SDS case ($t_{f}/\tau_{d}=0.4~{}t_{c}/\tau_{d}$, dashed line). For the highest concentrations, the lifetime follows this trend again, showing that the drainage has reached an evaporation dominated regime. The longest lifetimes obtained are similar for the cases with pure water or either salts and equal to about one minute. The case with NaCl and the largest amount of surfactant is noticeably above the rest of the data points, following the same linear trend but with a prefactor close to 0.6. It should be noted that in both cases this regime corresponds to an almost saturated isotherm (with a very low surface tension that does not vary with surface compression), similar to what would be obtained for a solution above the CMC. Finally, in the case of sea salt, the lifetime data departs from the trend at higher $t_{c}$ (high relative humidity or low wind speed) even at the highest SDS concentration, and increases slowly with $t_{c}$.
The behavior observed for surface bubbles with salts can qualitatively be replicated numerically by integrating Eq. (5) until the thickness reaches a finite thickness $h_{c}$. Results of this integration for various $h_{c}$ are shown in Fig. 7 in shades of orange, with the same numerical prefactors as in the case of SDS only ($\alpha_{d}=0.5$ and $\alpha_{e}=12.5$). With decreasing $h_{c}$ we see a transition from a regime where $t_{f}<t_{c}$ and $t_{f}$ is independent of relative humidity or wind speed to the same regime described in the case of pure SDS where $t_{f}\sim t_{c}$. This behavior qualitatively matches the one observed for sea salt or NaCl, showing that taking into account the effect of wind and relative humidity through the evaporative rate, and the effect of the solution through an average bursting thickness is enough to describe most of the trends seen in this study. The model predicts that the average bursting thickness in the wind independent regime is $h_{c}/h_{0}\approx 0.05$ or $h_{c}\approx 0.6$ µm for sea salt and $h_{c}/h_{0}\approx 0.02$ or $h_{c}\approx 0.2$ µm for NaCl. These values are coherent with what would be expected from bubbles with a radius of 1.6 mm with sea salt, extrapolating the data from (Shaw and Deike,, 2024) to our smaller size. Both cases become evaporation dominated when $h_{c}$ is less than 100 nm. Note that we have made here the hypothesis that the drainage curves are not altered by the presence of salts: Shaw and Deike, (2024) have recently measured the drainage of surface bubbles in a large range of conditions and have shown in particular that the effect of sea salt on drainage is much smaller than other effects such as solution temperature for instance and that the evolution of the thickness is still well captured by $h\sim t^{-2/3}$ at early times.
@@ 5 Discussion
The data presented in this article illustrates two very different regimes of surface bubble lifetime. For pure water and SDS only, the lifetime is essentially independent of the SDS concentration (from about one thousandth of the CMC to above the CMC) and closely follows Eq. (10). When either NaCl or sea salt is added, the lifetime is greatly reduced for a given SDS concentration and shows a transition from being independent of wind speed at low SDS concentrations to being dominated by evaporation at high concentrations as in the case of pure water and SDS.
@@
We believe the difference between these two regimes lies in the presence of impurities in the solution, which are required to trigger a bursting event at a large thickness. A large thickness means here bubble caps thicker than 1 µm, in opposition to the very small thicknesses ($h\lesssim 0.1$ µm) attainable for long-lasting bubbles. A bubble of pure water and surfactant does not burst before evaporation becomes the dominant drainage mechanism and rapidly (linearly in time) removes material from the cap. The regime where evaporation is dominant has previously been identified as the upper limit on bubble lifetime (Poulain and Bourouiba,, 2018), the difference with our study is the use of the overflow to remove most co-surfactants from the interface and ensure a precise control of surface properties. Overflow seems an efficient way of also removing the small amount of impurities (other surface active compounds than the main surfactant) that may be present when using only deionized water and surfactants in addition to ensuring that to physicochemical properties of solution remain constant (see Sec. 3).
@@
Authors have previously hypothesized that surface tension inhomogeneities (for instance across the marginal regeneration patches) may be enough to trigger bursting of bubble cap (Lhuissier and Villermaux,, 2012). From our data it seems that with SDS only, these inhomogeneities are not enough to make the bubble burst. This fact is supported by that wind, which drastically alters the marginal regeneration patches and presumably any Marangoni flow that might be present within the bubble cap, does not seem to have an effect on lifetime or drainage beyond a change of the evaporative rate.
@@
Introducing salts changes the surface tension but also adds impurities to the solution. These impurities may then trigger large thickness bursting as theorized in Poulain et al., (2018) (Sec 6). The sea salt that we used contains many different salts but also other inorganic compounds and therefore more impurities than NaCl (which is 99% pure). This fact may help explain why for a given SDS concentration, the sea salt case always has a shorter lifetime than the NaCl case (5 versus 10 s in the lower SDS concentration limit). Finally, another clue towards the need for impurities to trigger bursting at a large thickness is a few experiments we performed with tap water instead of deionized water (therefore also containing impurities). The resulting mean lifetime was close to 5 s with a small concentration of SDS (10 µmol/L) in contrast to about 15 s with deionized water. We also performed experiments with Triton X-100, a non-ionic, slow adsorbing surfactant, which confirms the trends observed with SDS: for a given small concentration, the mean lifetime follows $t_{f}\sim t_{c}$ with surfactant alone, but it is a constant at around 5s when sea salt is added. These two test cases point towards impurities being the reason for the difference between the cases with and without salts, instead of an interaction specific to SDS and ions in solutions.
@@
Knowledge of the Gibbs isotherm of the solution considered is not enough to predict the expected lifetime of a surface bubble. Indeed, the amount of impurities cannot be directly inferred from the isotherm and therefore neither can the expected bursting thickness. A method of measuring the nature and concentration of impurities in a solution may therefore help in predicting the exact lifetime of a surface bubble.
@@
Acknowledgement This work was supported by NSF grant 2242512, NSF CAREER 1844932 and the Cooperative Institute for Modeling the Earth’s System at Princeton University to L.D.

### v3

@@ 4.3 Lifetime of solutions of pure surfactants
Figure 6: a Mean lifetime as a function of the Reynolds number ReRe, for various relative humidity and surfactant concentration. (b) Dimensionless lifetime as a function of $\tilde{j_{e}}^{-3/5}$. Markers show relative humidity and colors show SDS concentration. (red dashed line) Eq. (12): $t_{f,{\rm max}}=1.32\alpha_{e}^{-3/5}\alpha_{d}^{-2/5}\times\tilde{j_{e}}^{-3/5}$, with $\alpha_{d}=1.8$ and $\alpha_{e}=1.2$.
In this section, we analyze the lifetime $t_{f}$ of surface bubbles (from the instant the bubble reaches the surface until it spontaneously bursts) in a solution only containing SDS (no salts). Plotting our data for several wind speeds, relative humidities and SDS concentrations as a function of the Reynolds number (Fig. 6 (a)) shows a behavior similar to the one observed in Fig. 4 (b), for all SDS concentrations (color) and humidities (markers). The mean lifetime with little to no wind ranges from 30 to 60 seconds and decreases down to 10 seconds for all cases as the Reynolds number increases.
We can compute the maximal time $t_{f,{\rm max}}$ that one can expect a bubble to remain on the surface before bursting using Eq. (10). Following the derivation in Poulain and Bourouiba, (2018), in the limit when the bursting thickness $h_{c}$ is very small compared to the thickness at generation ($h_{c}\ll h_{0}$), this time is given by:
$t_{f,{\rm max}}=t_{c}\int_{\tilde{h_{c}}}^{\tilde{h_{0}}}\frac{\mathrm{d}\tilde{h}}{\alpha_{e}\tilde{j_{e}}+\alpha_{d}\tilde{h}^{5/2}}\approx t_{c}\left(\alpha_{e}\tilde{j_{e}}\right)^{-3/5}\alpha_{d}^{-2/5}\int_{0}^{\infty}\frac{\mathrm{d}\tilde{h}}{1+\tilde{h}^{5/2}},$ (11)
and the last integral can be evaluated to obtain:
$t_{f,{\rm max}}\approx 1.32\>t_{c}\left(\alpha_{e}\tilde{j_{e}}\right)^{-3/5}\alpha_{d}^{-2/5}.$ (12)
@@
We therefore plot our mean lifetime data (normalized by $t_{c}$) as a function of $j_{e}^{-3/5}$ in Fig. 6 (b) and the data indeed follows a linear trend. This graph reveals that most of the changes in mean lifetime $\left<t_{f}\right$> are captured using a single parameter that contains the effects of wind and humidity on the evaporative rate: $\tilde{j_{e}}=\left(\mathrm{Sh}_{N}+\mathrm{Sh}_{F}\right)\left(1-\mathcal{R}_{H}\right)$. This result confirms that our modelling of the Sherwood number in the large range of conditions (from quiescent to fully turbulent) is correct. In addition, it shows that there are no visible effects of the wind on bubble lifetime beyond altering the evaporative rate. In this case with no salts in the solution, the effects of the surfactant concentration are small and do not show a significant trend.
The maximal lifetime $t_{f,{\rm max}}$ computed from Eq. (12) (red dashed line) follows closely the data and in particular is in quantitative agreement with our experimental measurements of $\left<t_{f}\right$>. This last result is consistent with the fact that the distribution of lifetimes is narrow around its mean (see Sec. 5) such that the maximal and mean lifetime are of the same order of magnitude. We note that in some of the cases, the mean experimental lifetime is larger than the maximal expected lifetime. This discrepancy may be due to small experimental differences between the drainage and lifetime setups: the tank in which the drainage experiment is performed is smaller (3 cm in diameter instead of 7 cm) such that the bubble always remains in view of the high speed camera. The proximity of a larger bath decreases the evaporative rate and therefore increases the maximal possible lifetime. At very high concentrations (black markers) other possible effects include a reduced evaporation rate due to the extreme contamination of the interface (Takemura and Yabe,, 1999) or a slowed down drainage due to the rigidification of the interface.
@@
Poulain and Bourouiba, (2018) observed for bubbles in the large surfactant concentration limit, that the longest time a surface bubble can live is given by $t_{f,{\rm max}}$. With our setup in which we took great care in removing impurities (see Sec 3), we find that the average lifetime is of order $t_{f,{\rm max}}$ (instead of just the maximum). This result is valid for very large changes of SDS concentration, from values greater than the CMC down to just a few µmol/L (or $10^{-3}$ times the CMC). All of our lifetime data for SDS is well collapsed when plotted as $\left<t_{f}\right>/t_{c}=f(j_{e}^{-3/5})$. This result is particularly robust at high wind speeds (or low $j_{e}^{-3/5}$), which is coherent with our previous observations that mixing the air around the bath and the bubble improves the repeatability of lifetime experiment by preventing local fluctuations of the humidity field. At high $j_{e}^{-3/5}$, we observe lifetimes that can be larger than the trend, and in general more important fluctuations. These may be due to the difficulty in modeling the Sherwood number associated with natural convection.
@@ 4.4 Lifetime of solutions containing salts
The lifetime of bubbles in solutions of salt and surfactants is drastically different from that of bubbles in pure water and surfactants. The lifetime data is summed up in Fig. 7 for sea salt (a mixture of salts and inorganic compounds in their oceanic proportions) and NaCl only (35 g/L). The top row shows the dimensional mean lifetime as a function of $Re$ where we see that in cases with a higher SDS concentration (yellow) the behavior is similar to cases with SDS only, with a gradual decrease of the mean lifetime as $Re$ is increased. At low concentrations however ($c_{\textsc{SDS}}\lesssim 50$ µmol/L, blue), the lifetime is essentially independent of wind speed or relative humidity. In this regime, the lifetime of solutions with sea salt is about 5 s, and 10 s with NaCl. It is interesting to note that in both cases, in this limit the lifetime is again independent on the surfactant concentration within the range tested. Even though the concentrations in SDS are fairly small (less than 5% of the CMC), the combined presence of salts and surfactants results in a large variation of the surface tension (see Appendix A).
@@
Figure 7: (a) and (b) Mean lifetime as function of ReRe, for various relative humidity, surfactant concentration, for the sea salt and NaCl cases, respectively. (c) and (d) Dimensionless lifetime as a function of $\tilde{j_{e}}^{-3/5}$. The left column shows the dataset with sea salt and the right one with NaCl. The dashed lines represent $t_{f,{\rm max}}$ and are identical to Fig. 6 (c) , given by Eq. (12). The orange lines are obtained by integrating of Eq. (10) until the thickness reaches a critical thickness $h_{c}$. Markers show relative humidity and color shows SDS concentration.
The bottom row shows the same data as a function of $\tilde{j_{e}}^{-3/5}$ illustrating that in this case too, the dimensionless evaporative rate is an adequate parameter to combine the effects of changing wind speeds and relative humidities. We can now see in both cases a transition between a regime where bubble lifetime is much shorter than $t_{f,{\rm max}}$ (dashed black line) and independent of relative humidity or wind speed, and a regime again dominated by evaporation, with the lifetime given by Eq. (12), as in Fig. 6 for pure water and SDS. The transition occurs for SDS concentration of around 100 and 200 µmol/L, where the increase of lifetime with $\tilde{j_{e}}^{-3/5}$ is slower than predicted by Eq. (12) and the lifetime always remains inferior to the pure water and SDS case (dashed line). For the highest concentrations, the lifetime follows this trend again, showing that the drainage has reached an evaporation dominated regime.
@@
The case with NaCl and the largest amount of surfactant is noticeably above the rest of the data points, following the same linear trend but with a larger prefactor. It should be noted that in both cases this regime corresponds to an almost saturated isotherm (with a very low surface tension that does not vary with surface compression), similar to what would be obtained for a solution above the CMC. Finally, in the case of sea salt, the lifetime data departs from the trend at higher $\tilde{j_{e}}^{-3/5}$ (high relative humidity or low wind speed) even at the highest SDS concentration, and increases slowly with $\tilde{j_{e}}^{-3/5}$.
The transition observed for surface bubbles with salts can qualitatively be replicated numerically by integrating Eq. (10) until the thickness reaches a finite thickness $h_{c}$. We make here the hypothesis that the drainage curves obtained in Fig. 5 remain unchanged upon the addition of salts and therefore again use the same prefactors $\alpha_{d}=1.8$ and $\alpha_{e}=1.2$. This approximation is supported by results from Shaw and Deike, (2024) who showed that the effect of sea salt on drainage is much smaller than other effects (such as solution temperature for instance) and that the evolution of the thickness is still well captured by $h\sim t^{-2/3}$ at early times.
Results of this integration for various $h_{c}$ are shown in Fig. 7 in shades of orange. With decreasing $h_{c}$ we see a transition from a regime where the lifetime obtained is smaller than $t_{f,{\rm max}}$ and is independent of relative humidity or wind speed when $h_{c}\gtrsim h^{\star}$, to the evaporation dominated regime (obtained analytically) for $h_{c}\ll h^{\star}$ where the lifetime is given by $t_{f,{\rm max}}$. This transition qualitatively matches the one observed for sea salt or NaCl, showing that taking into account the effect of wind and relative humidity through the evaporative rate, and the effect of the solution through an average bursting thickness is enough to describe most of the mean lifetime trends seen in this study.
The model predicts that the average bursting thickness in the wind independent regime is $h_{c}/h^{\star}\approx 1$ or $h_{c}\approx 0.6$ µm for sea salt and $h_{c}/h^{\star}\approx 0.5$ or $h_{c}\approx 0.3$ µm for NaCl. Both cases become evaporation dominated when $h_{c}$ is less than 100 nm. Values at rupture between $h_{c}\approx 0.3$ to $0.6$ µm are coherent with previous measurements on the drainage and rupture of bubbles in the presence of salts (Shaw and Deike,, 2024).
5 Lifetime probability distribution
Figure 8: Distributions of lifetimes in various cases. (a) Dimensional distributions of lifetime for various evaporation rates (color) with a solution of 10 µmol/L of SDS. (b) Same data, but each pdf is normalized by the mean lifetime, (black line with markers) Combined pdf of all cases. (c) Combined pdf for selected solutions, legend indicates $c_{\textsc{SDS}}$ in µmol/L and the type of salt. (d) Comparisons of the combined pdf with and without overflow. In (b) and (c) the dashed black line is a Weibull distribution of shape $4/3$ and scale $\Gamma(7/4)^{-1}$.
[…]
Using this approach, we can clearly see different bursting mechanisms in action for different solutions. In the case of deionized water with only SDS and no surfactants, the distributions of lifetime are extremely narrow around the mean, regardless of the concentration (Fig. 8, (c) green curves). This regime, as was described above, corresponds to a mean lifetime set by the evaporative rate, with a low likelihood of a bursting event before the time given by $t_{f,{\rm max}}$ (Eq. 12). We can also notice here that the width of the distribution is similar for low and high concentrations.
When salts are present (blue and orange curves), two regimes need to be distinguished: at low SDS concentrations, the addition of salt widens the distribution, but at high contamination, all cases are peaked with a similar shape as in the case without salt. Sea salt has a larger effect on the distribution than NaCl only. These observations are fully consistent with the mean lifetime approach of the rest of the study: the case without salt bursts at a time given by the evaporative rate, regardless of concentration. This results in a very narrow pdf around $t_{f,{\rm max}}$. For the cases with salts, at low SDS concentration the lifetime is independent of the evaporative rate and the bubbles can burst for a relatively wide range of lifetimes. At high concentration however the lifetime is dominated by evaporation, as seen from the mean lifetime, and the pdf is therefore also peaked.
In order to describe the distributions of bubble lifetime Lhuissier and Villermaux, (2012) have derived the following expression:
$\mathbb{P}\left(\frac{t_{f}}{\left<t_{f}\right>}\right)=\frac{4}{3}\Gamma\left(7/4\right)^{4/3}\left(\frac{t_{f}}{\left<t_{f}\right>}\right)^{1/3}\exp\left(-\left(\Gamma\left(7/4\right)\frac{t_{f}}{\left<t_{f}\right>}\right)^{4/3}\right),$ (13)
which is a Weibull distribution of shape $4/3$ and scale $\Gamma(7/4)^{-1}$, with $\Gamma$ the gamma function. This parameter-less distribution has been successfully used to describe lifetime data in several studies (Poulain et al.,, 2018; Miguet et al., 2020a, ; Lorenceau and Rouyer,, 2020). However, when plotted along our data (dashed black curve), we see that it does not describe any of the cases presented in this study: this distribution is extremely wide, and even in the cases with low SDS concentration and salt, the experimental distribution of lifetime is narrower around the mean than the proposed distribution Eq. (13).
To understand this discrepancy, we plot in panel (d) the distributions obtained when comparing similar cases with (blue) and without (green) overflow. At high SDS concentration (about the CMC), all cases are evaporation-dominated and therefore have a similarly narrow distribution. At lower concentrations however, the case with overflow remains very narrow while the case without overflow perfectly agrees with Eq. (13). The interpretation of the change in lifetime distribution is discussed below.
@@ 6 Discussion
The data presented in this article illustrates two very different regimes of surface bubble lifetime. For pure water and SDS only, the mean lifetime is essentially independent of the SDS concentration (from about one thousandth of the CMC to above the CMC) and closely follows the maximal possible lifetime given by Eq. (12), with a peaked distribution. When either NaCl or sea salt is added, the mean lifetime is greatly reduced for a given SDS concentration and the distribution is wider. We then see a transition from being independent of wind speed at low SDS concentrations to being dominated by evaporation at high concentrations as in the case of pure water and SDS.
@@
We believe the difference between these two regimes lies in the presence of impurities in the solution, which are required to trigger a bursting event at a large thickness. A large thickness means here bubble caps thicker than 1 µm, in opposition to the very small thicknesses ($h\lesssim 0.1$ µm) attainable for long-lasting bubbles. A bubble of pure water and surfactant does not burst before evaporation becomes the dominant drainage mechanism and rapidly (linearly in time) removes material from the cap. The regime where evaporation is dominant has previously been identified as the upper limit on bubble lifetime (Poulain and Bourouiba,, 2018), the difference with our study is the use of the overflow to remove most co-surfactants from the interface and ensure a precise control of surface properties. Overflow seems an efficient way of also removing the small amount of impurities (other surface active compounds than the main surfactant) that may be present when using only deionized water and surfactants in addition to ensuring that the physicochemical properties of solution remain constant (see Sec. 3).
@@
Authors have previously hypothesized that surface tension inhomogeneities (for instance across the marginal regeneration patches) may be enough to trigger bursting of bubble cap (Lhuissier and Villermaux,, 2012). From our data it seems that with SDS only, these inhomogeneities are not enough to make the bubble burst. This fact is supported by the observation that wind, which drastically alters the marginal regeneration patches and presumably any Marangoni flow that might be present within the bubble cap, does not seem to have an effect on lifetime or drainage beyond a change of the evaporative rate.
@@
Introducing salts changes the surface tension but also adds impurities to the solution. These impurities may then trigger large thickness bursting as theorized in Poulain et al., (2018) (Sec 6). The sea salt that we used contains many different salts but also other inorganic compounds and therefore more impurities than NaCl (which is 99% pure). This fact may help explain why for a given SDS concentration, the sea salt case always has a shorter lifetime than the NaCl case (5 versus 10 s in the lower SDS concentration limit) and a broader distribution.
Finally, another clue towards the need for impurities to trigger bursting at a large thickness is a few experiments we performed with tap water instead of deionized water (therefore also containing impurities). The resulting mean lifetime was close to 4 s with a small concentration of SDS (10 µmol/L) in contrast to about 15 s with deionized water. We also performed experiments with Triton X-100, a non-ionic, slow adsorbing surfactant, which confirms the trends observed with SDS: for a given small concentration, the mean lifetime follows $\left<t_{f}\right>\sim t_{f,{\rm max}}$ with surfactant alone, but it is a constant at around 4 s when sea salt is added. These two test cases (shown in App. B) point towards impurities being the reason for the difference between the cases with and without salts, instead of an electrostatic interaction specific to SDS and ions in solutions.
The observations on bubble lifetime distribution are also fully consistent with the interpretation that the presence of impurities in solution is responsible for the large thickness bursting regime. The cases with little impurities (with overflow and no salt) display a very peaked distribution around the mean, with a similar relative standard deviation across all cases. Introducing impurities widens the distribution by allowing the bubble to burst earlier than $t_{f,{\rm max}}$. Finally, the expression proposed by Lhuissier and Villermaux, (2012) can be though of as the limit at a large amount of impurities, where the cases without overflow aggregate. At very high surfactant concentration, the bubble cap is stable enough so that effect of impurities is less, and we recover peaked distributions and mean lifetimes given by the evaporative rate.
@@
Knowledge of the surface tension isotherm of the solution considered is not enough to predict the expected lifetime of a surface bubble. Indeed, the amount of impurities cannot be directly inferred from the isotherm and therefore neither can the expected bursting thickness. A method of measuring the nature and concentration of impurities in a solution may therefore help in predicting the exact lifetime of a surface bubble.
@@
Acknowledgement We thank an anonymous reviewer for their insightful comments that have significantly improved the theoretical derivation in this manuscript. This work was supported by NSF grant 2242512, NSF CAREER 1844932 and the Cooperative Institute for Modeling the Earth’s System at Princeton University to L.D.